# Topology and Spectrum in Measurement-Induced Phase Transitions

### v1

@@ Acknowledgements.
Acknowledgements— We thank Xhek Turkeshi, Henning Schomerus, Alessandro Romito, Amos Chan, Keiji Saito, Ken Shiozaki, and Takahiro Morimoto for valuable discussions and comments. We thank Zhenyu Xiao and Kohei Kawabata for coordinating the submission of this work and Ref. [72]. H.O. is supported by RIKEN Junior Research Associate Program. K. M. and R.H. are supported by JST ERATO Grant Number JPMJER2302, Japan. K.M. is supported by JSPS KAKENHI Grant No. JP23K13037. R.H. is supported by JSPS KAKENHI Grant No. JP24K16982. Y.F. is supported by JSPS KAKENHI Grants No. JP20K14402 and No. JP24K06897. Part of the computation has been performed using the facilities of the Supercomputer Center, the Institute for Solid State Physics, the University of Tokyo.

### v2

@@ Acknowledgements.
Acknowledgments— We thank Xhek Turkeshi, Henning Schomerus, Alessandro Romito, Amos Chan, Keiji Saito, Ken Shiozaki, and Takahiro Morimoto for valuable discussions and comments. We thank Zhenyu Xiao and Kohei Kawabata for coordinating the submission of this work and Ref. [72]. H.O. is supported by RIKEN Junior Research Associate Program. K. M. and R.H. are supported by JST ERATO Grant Number JPMJER2302, Japan. K.M. is supported by JSPS KAKENHI Grant No. JP23K13037. R.H. is supported by JSPS KAKENHI Grant No. JP24K16982. Y.F. is supported by JSPS KAKENHI Grants No. JP20K14402 and No. JP24K06897. Part of the computation has been performed using the facilities of the Supercomputer Center, the Institute for Solid State Physics, the University of Tokyo.